OpenStudy (anonymous):

How do you factor the polynomial x2-18x+45

OpenStudy (anonymous):

x2 - 18x + 45 multiply / \ V V ( x - 3 ) ( x - 15 ) x2 - 18x + 45 multiply / \ V V ( x - 3 ) ( x - 15 ) ^ ^ \ / add x2 - 18x + 45 ( x - 3 ) ( x - 15 ) this is in its factored form if you were to solve it ( x - 3 ) ( x - 15 ) = 0 ( x - 3 ) = 0 ( x - 15 ) = 0 x = { 3, 15} both forms of the answer should be correct in value x = 3 x = 15